# VerifyThisBench: Generating Code, Specifications, and Proofs All at Once

### v1

@@ Abstract
Large language models (LLMs) have demonstrated remarkable progress in code generation, but many existing benchmarks are approaching saturation and offer little guarantee on the trustworthiness of the generated programs, offering limited insight into deeper reasoning capabilities. We introduce VerifyThisBench, a new benchmark designed to evaluate LLMs on end-to-end program verification tasks that require interpreting natural language problem descriptions, formulating formal specifications, generating code, and constructing correctness proofs. Our evaluation reveals that even state-of-the-art (SOTA) models, such as o3-mini, achieve a pass rate of less than 4%, with many outputs failing to compile. To reduce task complexity, we further propose VerifyThisBenchXS, a variant in which partial implementations or proofs are provided. We systematically assess SOTA models on both benchmarks, uncovering key strengths and limitations in their formal reasoning and verification capabilities.
@@ 1 Introduction
Large Language Models (LLMs) have unequivocally revolutionized the landscape of automated code generation. Models like GPT-4o [33], Google Gemini [15], Anthropic Claude [1], and GitHub Copilot [14] excel at generating functional code snippets and translating between languages. These capabilities are now integrated into AI-powered IDEs, such as Cursor AI, to support large-scale software development. This proficiency had led to increasing needs on established benchmarks, as early as HumanEval [6] and MBPP [3], to reflect the capability of each LLM tool. However, this rapid progress raises critical questions about the trustworthiness and reliability of the generated artifacts. Many existing benchmarks, while useful for gauging functional correctness through test suites, are approaching saturation [21, 13, 16, 45] and inherently offer limited guarantees regarding the deeper aspects of program correctness. Test cases, by their nature, can demonstrate the presence of bugs, but cannot prove their absence [11], leaving a significant gap in assessing the formal robustness and true reasoning capabilities of these powerful models.
@@
Reliable software must go beyond passing tests to be trustworthy, precisely follow specifications, and even self-validate. Formal verification offers the most rigorous approach to achieving these guarantees. This paradigm involves providing mathematical proof that a program adheres to a formal specification, thereby guaranteeing critical properties such as functional correctness, liveness (ensuring the program eventually does something good), and safety (ensuring the program never does something bad) [18]. Modern program verification infrastructures, such as Dafny [27], Frama-C [22], Verus [26], Isabelle/HOL [32], and Lean [8], coupled with powerful automated theorem provers and SMT solvers like Z3 [7] and CVC5 [4], have significantly streamlined the process of writing and checking such verified software. These tools allow developers to express complex specifications and then automatically or semi-automatically verify that the implementation meets these specifications.
@@
Although researchers have developed multiple benchmarks to assess LLMs on formal verification subtasks [20, 5, 35, 12], none evaluate end-to-end program verification starting from a natural-language description. Instead, existing suites either require verifying or synthesizing small programs against a formal specification or focus on aiding proof completion by suggesting individual verification steps. Consequently, even though state-of-the-art LLMs have been reported to solve up to 97.8% of these benchmark tasks [44], those numbers do not reflect their true capability for end-to-end program verification.
@@
To bridge this gap and rigorously evaluate the capabilities of LLMs in this demanding domain, we introduce VerifyThisBench, a novel benchmark designed to assess end-to-end program verification, as shown in Figure 1. Inspired by the annual VerifyThis Challenge [43], where human contestants devise implementations and accompanying formal proofs in verification-aware languages, VerifyThisBench tasks LLMs with interpreting natural language problem descriptions, formulating formal specifications, generating the corresponding code, and constructing machine-checkable correctness proofs—all at once.
@@
Our evaluation using VerifyThisBench reveals that even state-of-the-art (SOTA) models, such as o3-mini, achieve a zero-shot pass rate of 3.62% on this end-to-end task, with a significant number of outputs failing even to compile, and only reach a pass rate of 9.37% after five rounds of feedback. These results underscore the profound challenge this domain presents. To dissect these challenges further and explore capabilities in a more guided setting, we also propose VerifyThisBenchXS, a variant where partial implementations or proofs are provided, and the LLM’s task is to complete the missing components.
@@
  • SOTA LLM Evaluation: We conduct a systematic evaluation of nine SOTA LLMs on both benchmarks, revealing current capabilities and significant limitations. Our analysis includes performance breakdowns across tools, attempt-based comparisons, model-specific strengths, self-assessed coherence, and the impact of partial guidance, providing a comprehensive understanding of model behavior in formal verification tasks.
@@ 2.1 Unverified Code Synthesis Benchmarks
Recent benchmarks for code generation include APPS [17], HumanEval [6], MBPP [3], CodeContests [28], DS-1000[25], SWEBench [19], and EvalPlus [29], among others. These benchmarks present programming tasks, often sourced from online competitions or community platforms, and evaluate models based on whether generated solutions pass a set of input-output test cases. While effective for measuring functional correctness, they do not involve formal specifications, proofs, or specification synthesis.
In contrast, VerifyThisBench requires models to go beyond functional testing: they must extract a formal specification from a natural-language description, generate code in a verification-aware language, and produce a proof that passes formal verification. This makes VerifyThisBench a substantially more rigorous and comprehensive benchmark than traditional synthesis tasks.
@@ 2.2 Program Verification Benchmarks
Benchmarks in formal verification include SV-COMP [37], SyGuS [38], and Code2Inv [36]. SV-COMP and Code2Inv focus solely on verification task with no code generation involved. Specifically, the former contains large-scale C benchmarks with fixed safety properties and the latter targets invariant generation over small C-style programs. SyGuS focuses on constraint-based synthesis.
More recent efforts like DafnyBench  [30] and VerusBench  [46] collect verified programs in Dafny and Verus respectively, primarily to train and evaluate ML-based tools for aiding in proof completion and suggesting verification steps, rather than end-to-end program generation from natural language.
@@ 2.3 Formal Methods in Software Verification: A Primer
    Annotation: Developers write code in a verification-aware language (e.g., Dafny [27], Frama-C [22], Verus [26] and annotate it with formal specifications and proof hints, including pre-conditions, post-conditions, assertions, and loop invariants.
@@
Successfully generating code within this paradigm, as targeted by our VerifyThisBench benchmark, requires an LLM not only to produce the algorithmic implementation but also to understand, formulate, and correctly express these intricate formal specifications and proof structures that enable automated verification.
@@ 3 VerifyThisBench Benchmark
VerifyThisBench is inspired by the annual VerifyThis Challenges [43], a competition where participants are tasked with formalizing specifications, implementing solutions, and verifying that the implementations meet the specification. Each challenge is designed to be completed within a 90-minute session and varies in difficulty. Submissions are evaluated based on correctness, completeness, and additional quality criteria such as elegance and the degree of automation. Similarly, in VerifyThisBench, the task is to interpret natural language problem descriptions and implement code and write proofs.
@@ 3.1 Benchmark Construction
We collected challenges from each annual competition between 2011 and 2024, documenting their descriptions, pseudo-code, and associated tasks. Tasks are categorized as either implementation (completing an algorithm from pseudo-code) or verification (proving a model or implementation correct against a specification). All tasks are described in natural language. In total, the dataset includes 41 challenges and 154 tasks. The dataset is available at [10].
@@ 3.2 Environment
To facilitate evaluation, we provide a unified environment supporting seven verification tools. Five of them, Dafny [27], Why3 [42], VeriFast [41], VerCors [40], and Frama-C [22], are widely used in past VerifyThis competitions. To broaden tool diversity, we additionally include Verus [26] and CBMC [23]. Tool versions and brief descriptions can be found in Appendix D.
@@ 3.3 Features of VerifyThisBench
End-to-end verification tasks with natural language problem descriptions: All tasks start with informal, natural language prompts (often with pseudo-code). Models must interpret the intent and formalize it into precise logical specifications. They are required to generate specifications, implementations, and formal proofs in a verification-aware language, ensuring the code passes machine-checkable verification. Example challenge and solution can be found in Appendix E.
Graded difficulty and multi-step challenges: Challenges are drawn from the VerifyThis competition and span a range of difficulties. Many include sequential subtasks, allowing fine-grained assessment of model capability.
Tool diversity: Multiple tools are provided. Models must conform to the syntax and semantics of real-world verification frameworks.
@@ 3.4 Relaxation
We observe that most language models fail to generate compilable code when targeting specific formal verification tools. This is often due to the syntactic complexity and precise annotations required by these tools. To decrease the level of difficulty and better assess LLM capabilities under more supportive conditions, we construct a set of relaxed problems derived from past human-written solutions. Specifically, we define three forms of relaxation:
  • Code relaxation: We provide only the function specifications, omitting both the implementation and the proof annotations.
  • Specification relaxation: We provide the implementation and its proof, but remove the function specifications.
  • Proof relaxation: We provide specifications and implementations, but remove loop invariants and other auxiliary annotations needed for verification.
To further diversify the difficulty spectrum, we vary the extent of relaxation. In some instances, we remove all relevant components (e.g., entire specs or proofs), while in others, we retain partial elements or include complete examples as guidance. This enables a more graded evaluation of LLM performance across varying levels of verification support. In total, we create a set of 481 tasks. Specifically, there are 195 fill-implementation task, 90 fill-proof/invariants tasks, and 196 fill-specification tasks. Table 6 in Appendix A shows the statistics of VerifyThisBenchXS. As there are no prior solutions in CBMC and Verus, no tasks were created and no results reported for these tools in the relaxed setting.
@@ 4 Experiment Results
4.1 Evaluation Setup and Metrics
We evaluate a diverse set of state-of-the-art (SOTA) language models, including both proprietary and open-source systems. The models include representatives from the OpenAI family (GPT-4o, GPT-4omini, o3-mini, o4-mini) [34], Anthropic (Claude-3.7-Sonnet) [2], Google (Gemini-2.5-Flash), DeepSeek (Deepseek-chat-v3) [9], Meta (Llama3.3-70B-Instruct) [31] and Alibaba (Qwen-2.5-72B-Instruct) [39]. This selection enables a comprehensive comparison across different model architectures and training paradigms. Model versions are provided in appendix C.
4.2 Experiment Design
For both VerifyThisBench and VerifyThisBenchXS, we conduct experiments with iterative refinement based on tool-generated error messages. To evaluate correctness, we pass the generated code to the target verification tool and check whether it compiles and verifies successfully. A task is marked as pass if no errors are returned.
In addition to correctness checking, we introduce a coherence check as a relaxed evaluation metric. Here, the model self-assesses whether its generated code semantically aligns with the original problem intent—an aspect difficult to verify automatically. This helps evaluate how well the specification matches the task description and provides insight into the model’s confidence in its output.
Each task is attempted five times per model. The first attempt uses only the task prompt; the next four incorporate feedback from previous errors. During refinement, the model has access to the full history of its prior attempts and corresponding feedback for the current task, enabling iterative improvement.
In VerifyThisBench, tasks of a challenge are completed sequentially. Only the final attempt from the previous task is carried over to the next, preserving essential context while keeping prompts concise. In contrast, VerifyThisBenchXS tasks have isolated contexts and are completed independently, with no progress carried over between tasks.
To ensure fairness, we use the same prompt across all models and set the temperature to 0.7 when applicable. A timeout of 1 minute is enforced for all experiments on the verifier. The experiments were conducted on a machine with an Intel i7-1360P CPU and 16GB of RAM.
@@ 4.3 Overall Pass Rate
Table 1: Overall Pass Rate On VerifyThisBench
CBMC zero-shot 8.44% 7.14% 8.44% 1.30% 6.49% 1.95% 7.14% 0.65% 1.30%
refinement 20.13% 19.48% 25.32% 15.58% 22.08% 14.94% 20.13% 22.08% 3.25%
Dafny zero-shot 0 0 4.55% 1.95% 3.90% 0 0 1.30% 0
refinement 1.30% 0.65% 9.74% 10.39% 11.04% 1.30% 2.60% 2.60% 0.65%
Frama-C zero-shot 0 0.65% 0 0 3.90% 0.65% 0 1.95% 0
refinement 7.14% 1.95% 2.60% 3.25% 11.04% 8.44% 0.65% 3.25% 0
VerCors zero-shot 0 1.30% 1.30% 1.95% 0 5.84% 8.44% 1.95% 0
refinement 1.95% 1.30% 1.95% 5.19% 1.30% 16.88% 11.69% 4.55% 3.25%
VeriFast zero-shot 0 0 0 0 0 0 0 0 0
refinement 0 0 0 2.60% 0 0 0.65% 0.65% 0
Verus zero-shot 1.95% 6.49% 10.39% 0.65% 0.65% 0.65% 7.79% 0.65% 0.65%
refinement 12.99% 9.09% 21.43% 8.44% 1.30% 0.65% 17.53% 1.30% 0.65%
Why3 zero-shot 0 0 0.65% 0.65% 1.30% 1.30% 0 0.65% 0
refinement 0 0 4.55% 10.39% 9.09% 5.84% 1.95% 1.95% 0
Overall zero-shot 1.48% 2.23% 3.62% 0.93% 2.32% 1.48% 3.34% 1.02% 0.28%
refinement 6.22% 4.64% 9.37% 7.98% 7.98% 6.86% 7.88% 5.19% 1.11%
Improvement 4.73% 2.41% 5.75% 7.05% 5.66% 5.38% 4.55% 4.17% 0.83%
Table 2: Overall Pass Rate On VerifyThisBenchXS
Dafny zero-shot 0 1.35% 2.70% 1.35% 1.35% 1.35% 0 2.70% 1.35%
refinement 17.57% 9.46% 31.08% 37.84% 41.89% 17.57% 8.11% 21.62% 8.11%
Frama-C zero-shot 0 0 1.85% 0 1.85% 1.85% 0 1.85% 0
refinement 0 0 5.56% 18.52% 9.26% 1.85% 0 5.56% 0
VerCors zero-shot 0 0 0 0 0 0 0 0 0
refinement 0 0 0 7.69% 0 0 3.85% 0 0
VeriFast zero-shot 7.58% 4.55% 3.03% 6.06% 6.06% 0 4.55% 12.12% 3.03%
refinement 12.12% 6.06% 4.55% 10.61% 27.27% 0 9.09% 13.64% 6.06%
Why3 zero-shot 0 0 0 0.38% 0.38% 0.77% 0 0.38% 0
refinement 7.66% 2.30% 0.77% 8.81% 3.45% 1.15% 15.71% 18.77% 1.15%
Overall zero-shot 1.04% 0.83% 1.04% 1.25% 1.46% 0.83% 0.62% 2.49% 0.62%
refinement 8.52% 3.53% 6.44% 14.55% 13.10% 3.53% 11.23% 16.01% 2.70%
Improvement 7.48% 2.70% 5.40% 13.30% 11.64% 2.70% 10.61% 13.52% 2.08%
Table 1 presents the performance of the SOTA models on VerifyThisBench. For each verification tool, we report pass rates for the initial zero-shot attempt and after four additional refinement attempts using feedback.
@@
Table 2 shows the results on VerifyThisBenchXS. Similarly, at the first attempt, absolute numbers remain low (less than 3%) for all models. At the fifth iteration, Deepseek tops the competition with 16.01%, followed closely by o4-mini (14.55%), Claude (13.10%), and Llama (11.23%). Feedback leads to substantial improvement for most models, achieving relative gains of over 10%.
@@
Key Insights: Average pass rates remain low at 10% on VerifyThisBench and 16% on VerifyThisBenchXS, revealing the challenges formal verification poses even to SOTA LLMs. All models improve with feedback.

### v2

@@ Abstract
Large language models (LLMs) have demonstrated remarkable progress in code generation, but many existing benchmarks are approaching saturation and offer little guarantee on the trustworthiness of the generated programs. To improve visibility into model reasoning on formal correctness, we introduce VerifyThisBench, a new benchmark that evaluates end‑to‑end program verification from natural language descriptions: models must (i) extract formal specifications, (ii) implement in a verification‑aware language, and (iii) construct machine‑checkable proofs. Our evaluation reveals that even state-of-the-art (SOTA) models, such as o3-mini, achieve a pass rate of less than 4%, with many outputs failing to compile. To isolate sources of difficulty, we further propose VerifyThisBenchXS, a relaxed variant in which partial implementations or proofs are provided. Across nine models and seven verification tools on both benchmarks, we observe consistent gains with feedback‑driven refinement, but overall pass rates remain low, underscoring substantial gaps in formal reasoning. We release the benchmark and the unified evaluation environment to catalyze the verification capabilities for future models.
@@ 1 Introduction
Large language models (LLMs) have unequivocally revolutionized the landscape of automated code generation. Models like ([34]) GPT-4o, ([16]) Gemini, ([1]) Claude, and ([15]) Copilot excel at generating functional code snippets and translating between languages. These capabilities are now integrated into AI-powered IDEs, such as Cursor and Visual Studio, to support large-scale software development. This proficiency had led to increasing needs on established benchmarks, as early as HumanEval [6] and MBPP [3], to reflect the capability of each LLM tool. However, this rapid progress raises critical questions about the trustworthiness and reliability of the generated artifacts. Many existing benchmarks, while useful for gauging functional correctness through test suites, are approaching saturation ([22, 14, 17, 47]) and inherently offer limited guarantees regarding the deeper aspects of program correctness. Test cases, by their nature, can demonstrate the presence of bugs, but cannot prove their absence [12], leaving a significant gap in assessing the formal robustness and true reasoning capabilities of these powerful models.
@@
Reliable software must go beyond passing tests to be trustworthy, precisely follow specifications, and even self-validate. Formal verification offers the most rigorous approach to achieving these guarantees. This paradigm involves providing machine-checked mathematical proofs to show that a program adheres to its formal specification, thereby guaranteeing critical properties such as functional correctness, liveness (ensuring the program eventually does something good), and safety (ensuring the program never does something bad) [19]. Modern program verification infrastructures, such as Dafny [28], Frama-C [23], Verus [27], Isabelle/HOL [33], and Lean [8], coupled with powerful automated theorem provers and SMT solvers like Z3 [7] and CVC5 [4], have significantly streamlined the process of writing and checking such verified software. These tools allow developers to express complex specifications and then automatically or semi-automatically verify that the implementation meets these specifications.
@@
Although researchers have developed multiple benchmarks to assess LLMs on formal verification subtasks [21, 5, 36, 13], none evaluates end-to-end program verification solely from natural-language inputs. Instead, existing suites either require verifying or synthesizing small programs against a given formal specification, or focus on aiding proof completion by suggesting individual verification steps. Consequently, even though state-of-the-art LLMs have been reported to solve up to 97.8% of these benchmark tasks [46], those numbers do not reflect their true capability for end-to-end program verification.
@@
To bridge this gap and rigorously evaluate the capabilities of LLMs in this demanding domain, we introduce VerifyThisBench, a novel benchmark designed to assess end-to-end program verification, as shown in Figure 1. Inspired by the annual ([44]) where human contestants devise implementations and accompanying formal proofs in verification-aware languages, VerifyThisBench tasks LLMs with interpreting natural language problem descriptions, formulating formal specifications, generating the corresponding code, and constructing machine-checkable correctness proofs – all at once, to produce compiled and verified artifacts. While recent efforts  [49, 41] also benchmark LLMs on end-to-end verification tasks in Lean, our work differs by building on the long-standing VerifyThis Challenge, offering multi-framework coverage, research-grade tasks, and competition-vetted difficulty, with solution lengths up to 648 lines compared to a maximum of 225 in prior work.
@@
Our evaluation using VerifyThisBench reveals that even state-of-the-art (SOTA) models, such as o3-mini, achieve a zero-shot pass rate of 3.62% on this end-to-end task, with a significant number of outputs failing even to compile, and only reach a pass rate of 9.37% after five rounds of feedback. These results underscore the profound challenge this domain presents. To dissect these challenges further and explore capabilities in a more guided setting, we also propose VerifyThisBenchXS, a variant where partial specification, implementation code, or proofs are provided, and the LLM is tasked to complete the missing components. In this setting, o3-mini achieves 2.24% in zero-shot attempt and 8.28% after refinement.
@@
  • SOTA LLM Evaluation: We conduct a systematic evaluation of nine SOTA LLMs on both benchmarks, revealing current capabilities and significant limitations.
@@ 2.1 Unverified Code Synthesis Benchmarks
Recent benchmarks for code generation include APPS [18], HumanEval [6], MBPP [3], CodeContests [29], DS-1000[26], SWEBench [20], and EvalPlus [30], among others. These benchmarks present programming tasks, often sourced from online competitions or community platforms, and evaluate models based on whether generated solutions pass a set of input-output test cases. While effective in emulating daily software development, they do not involve formal verification.
In contrast, VerifyThisBench requires models to go beyond functional testing: they must formalize natural language intents into specifications, generate code in verification-aware languages, and produce proofs that pass a formal logic verifier. This makes VerifyThisBench a substantially more rigorous and comprehensive benchmark than traditional code synthesis tasks.
@@ 2.2 Program Verification Benchmarks
Benchmarks built in the context of formal verification include SV-COMP [39], SyGuS [40], and Code2Inv [38]. SV-COMP and Code2Inv focus solely on verification tasks that do not require implementation generation. For more contexts, the former contains large-scale C/Java benchmarks verifying fixed safety properties, and the latter targets loop-invariant generation over small C-style programs. SyGuS focuses on constraint-based synthesis.
More recent efforts like DafnyBench  [31] and VerusBench  [48] collect verified programs in Dafny and Verus respectively, primarily to train and evaluate ML-based tools in aiding proof completion and suggesting verification steps, rather than end-to-end program generation from natural language.
@@
2.3 End-to-end Verification Benchmarks
Parallel work includes Verina [49] (189 tasks) and Clever [41] (161 tasks), exploring end-to-end verification in Lean, with sources translated from programming tasks in HumanEval, MBPP and Leetcode etc. VerifyThisBench differs in the source, scope and diversity, with 734 tasks derived from the VerifyThis competition series, which presents realistic, research-grade verification challenges across multiple domains. Rather than focusing on Lean, VerifyThisBench spans seven verification frameworks across multiple programming languages, including Verus [27] and Frama-C [23] that are established for production codebases. Moreover, VerifyThisBench includes tasks that require reasoning about memory safety, concurrency, and complex data structures beyond arrays and trees. See Appendix E for detailed comparison of description lengths and solution sizes.
@@ 2.4 Formal Methods in Software Verification: A Primer
    Annotation: Developers write code in a verification-aware language (e.g., Dafny [28], Frama-C, Verus and annotate it with formal specifications and proof hints, including pre-conditions, post-conditions, assertions, and loop invariants.
@@
Successfully generating code within this paradigm, as targeted by our VerifyThisBench, requires an LLM not only to produce the algorithmic implementation but also to understand, formulate, and correctly express intricate formal specifications and proof structures that enable automated verification.
@@ 3 VerifyThisBench Benchmark
VerifyThisBench is inspired by the annual VerifyThis Challenges [44], a competition where participants are tasked with formalizing specifications, implementing solutions, and verifying that the implementations meet the specification. We focus on this benchmark because it is a dedicated formal methods competition, designed not only to evaluate participants’ skills but also to assess the maturity of verification tools [11]. Each challenge is designed to be completed within a 90-minute session and varies in difficulty. Submissions are evaluated based on correctness, completeness, and additional quality criteria such as elegance and the degree of automation. Similarly, in VerifyThisBench, the task is to interpret natural language problem descriptions and implement code and write proofs.
@@ 3.1 Benchmark Construction
We collected challenges from the annual competition series between 2011 and 2024, each with natural-language descriptions (seldom include pseudo-code) and associated (one or more) tasks. Tasks are categorized as either implementation (completing an algorithm) or verification (proving a model or implementation correct against a specification). The resulting dataset includes 41 challenges and 154 tasks, with an example in Appendix H and detailed statistics in Appendix E. The dataset is available in supplementary material and will be made public after the anonymity period.
@@ 3.2 Environment
To facilitate evaluation, we provide a unified environment supporting seven verification tools. Five of them, Dafny [28], ([45]), ([43]), ([42]), and Frama-C [23], are widely used in past VerifyThis competitions. To broaden tool diversity, we additionally include Verus [27] and CBMC [24], covering Rust, C, and other imperative or deductive platforms. Tool versions and brief descriptions can be found in Appendix C.
@@ 3.3 Features of VerifyThisBench
End-to-end verification tasks with natural language problem descriptions: All tasks start with informal, natural language prompts (often with pseudo-code). Models must interpret the intent and formalize it into precise logical specifications. They are required to generate specifications, implementations, and formal proofs in a verification-aware language, ensuring the code passes machine-checkable verification. Example challenge and solution can be found in Appendix H.
Graded difficulty and multi-step challenges: Challenges are drawn from the VerifyThis competition and span a range of difficulties (see Appendix E.). Many include sequential subtasks, allowing fine-grained assessment of model capability on step-wise tasks.
Tool diversity: Multiple tools are provided and tested on. Models must conform to the syntax and semantics of real-world verification frameworks.
@@ 3.4 Relaxation
We observe that most language models fail to generate compilable code when targeting specific verification tools. This is often due to the syntactic complexity and precise annotations required by these tools. To isolate the sources of difficulty and better assess LLM capabilities under more supportive conditions, we construct a set of relaxed subtasks derived from past human-written solutions. Specifically, we define three forms of relaxation. In Code-Gen, we provide the function specifications, omitting both the implementation and the proof annotations. In Specification-Gen, we provide the implementation and its proof, but remove the function specifications. In Loop-Gen, we provide specifications and implementations, but remove loop invariants needed for verification.
In total, we create a set of 580 tasks. Specifically, there are 226 code-gen task, 121 loop-gen tasks, and 233 spec-gen tasks. Table 6 in Appendix A shows the statistics of VerifyThisBenchXS. Since no prior solutions exist for CBMC and Verus, and given notable community interest, we developed new Verus solutions to enrich the dataset; CBMC solutions remain unavailable and are therefore not included in the relaxed experiments.
@@ 4 Experiment Results
4.1 Model Setup
We evaluate a diverse set of SOTA language models, covering both proprietary and open-source systems. Representatives are selected from the ([35]) family (GPT-4o, GPT-4omini, o3-mini, o4-mini), ([2]) (Claude-3.7-Sonnet), Google (Gemini-2.5-Flash) [9], DeepSeek (Deepseek-chat-v3) [10], ([32]) (Llama3.3-70B-Instruct) and Alibaba (Qwen-2.5-72B-Instruct) [37]. This selection enables a comprehensive comparison across different model architectures and training paradigms. Model versions are provided in Appendix B.
4.2 Experiment Design and Metrics
For both VerifyThisBench and VerifyThisBenchXS, we conduct experiments with iterative refinement based on tool-generated error messages. To evaluate correctness, we pass the generated code to the target verification tool and check whether it compiles and verifies successfully. A task is marked as pass/succeed if no error is returned.
In addition to correctness checking, we introduce a coherence check as a relaxed evaluation metric. In this step, the model self-assesses whether its generated code semantically aligns with the original problem intent – an aspect difficult to verify automatically. This metric helps determine how well the specification matches the task description and provides insight into the model’s ability in auto‑formalization and symbol grounding.
Each task is attempted five times per model. The first attempt uses only the task prompt; the next four incorporate feedback from previous errors. During refinement, the model has access to the full history of its prior attempts and corresponding feedback for the current task, enabling iterative correction.
In VerifyThisBench, a challenge may have multi-stage tasks that are completed sequentially. Only the final attempt from the previous subtask is carried over to the next, preserving essential contexts while keeping the prompts concise. In contrast, VerifyThisBenchXS tasks have isolated contexts and are completed independently, with no progress carried over between tasks.
To ensure fairness, we use the same prompt (see Appendix D) across all models and set the temperature to 0.7 when applicable. Timeout of one minute is enforced for all experiments on the verifiers. The experiments were conducted on a machine with an Intel i7-1360P CPU and 16GB of RAM.
@@ 4.3 Overall Pass Rate
Table 1 presents the performance of the SOTA models on VerifyThisBench. For each verification tool, we report pass rates on the initial zero-shot attempt and after four additional refinement attempts using feedback.
@@
Table 1: Overall Pass Rate On VerifyThisBench
Attempt GPT4o GPT4o-mini o3-mini o4-mini Claude Gemini Llama Deepseek Qwen CBMC zero-shot 8.44% 7.14% 8.44% 1.30% 6.49% 1.95% 7.14% 0.65% 1.30% refinement 20.13% 19.48% 25.32% 15.58% 22.08% 14.94% 20.13% 22.08% 3.25% Dafny zero-shot 0 0 4.55% 1.95% 3.90% 0 0 1.30% 0 refinement 1.30% 0.65% 9.74% 10.39% 11.04% 1.30% 2.60% 2.60% 0.65% Frama-C zero-shot 0 0.65% 0 0 3.90% 0.65% 0 1.95% 0 refinement 7.14% 1.95% 2.60% 3.25% 11.04% 8.44% 0.65% 3.25% 0 VerCors zero-shot 0 1.30% 1.30% 1.95% 0 5.84% 8.44% 1.95% 0 refinement 1.95% 1.30% 1.95% 5.19% 1.30% 16.88% 11.69% 4.55% 3.25% VeriFast zero-shot 0 0 0 0 0 0 0 0 0 refinement 0 0 0 2.60% 0 0 0.65% 0.65% 0 Verus zero-shot 1.95% 6.49% 10.39% 0.65% 0.65% 0.65% 7.79% 0.65% 0.65% refinement 12.99% 9.09% 21.43% 8.44% 1.30% 0.65% 17.53% 1.30% 0.65% Why3 zero-shot 0 0 0.65% 0.65% 1.30% 1.30% 0 0.65% 0 refinement 0 0 4.55% 10.39% 9.09% 5.84% 1.95% 1.95% 0 Overall zero-shot 1.48% 2.23% 3.62% 0.93% 2.32% 1.48% 3.34% 1.02% 0.28% refinement 6.22% 4.64% 9.37% 7.98% 7.98% 6.86% 7.88% 5.19% 1.11% Improvement 4.73% 2.41% 5.75% 7.05% 5.66% 5.38% 4.55% 4.17% 0.83%
Table 2: Overall Pass Rate On VerifyThisBenchXS
Attempt GPT4o GPT4o-mini o3-mini o4-mini Claude Gemini Llama Deepseek Qwen Dafny zero-shot 0 1.35% 2.70% 1.35% 1.35% 1.35% 0 2.70% 1.35% refinement 17.57% 9.46% 31.08% 37.84% 41.89% 17.57% 8.11% 21.62% 8.11% Frama-C zero-shot 0 0 1.85% 0 1.85% 1.85% 0 1.85% 0 refinement 0 0 5.56% 18.52% 9.26% 1.85% 0 5.56% 0 VerCors zero-shot 0 0 0 0 0 0 0 0 0 refinement 0 0 0 7.69% 0 0 3.85% 0 0 VeriFast zero-shot 7.58% 4.55% 3.03% 6.06% 6.06% 0 4.55% 12.12% 3.03% refinement 12.12% 6.06% 4.55% 10.61% 27.27% 0 9.09% 13.64% 6.06% Verus zero-shot 7.07% 5.05% 8.08% 14.14% 14.14% 4.04% 3.03% 7.07% 5.05% refinement 15.15% 6.06% 17.17% 30.30% 30.30% 16.16% 13.13% 20.20% 7.07% Why3 zero-shot 0 0 0 0.38% 0.38% 0.77% 0 0.38% 0 refinement 7.66% 2.30% 0.77% 8.81% 3.45% 1.15% 15.71% 18.77% 1.15% Overall zero-shot 2.07% 1.55% 2.24% 3.45% 3.62% 1.38% 1.03% 3.28% 1.38% refinement 9.66% 3.97% 8.28% 17.24% 16.03% 5.69% 11.55% 16.72% 3.45% Improvement 7.59% 2.42% 6.04% 13.79% 12.41% 4.31% 10.52% 13.44% 2.07%
Table 2 shows the results on VerifyThisBenchXS. Similarly, at the first attempt, absolute numbers remain low (less than 4%) for all models. At the fifth iteration, o4-mini tops the competition with 17.24%, followed closely by Deepseek (16.72%), Claude (16.03%), and Llama (11.55%). Feedback leads to substantial improvement for most models, achieving relative gains of over 10%.
@@
Key Insights: Average pass rates for all evaluated models remain low at 10% on VerifyThisBench and 18% on VerifyThisBenchXS, revealing the challenges formal verification poses even to SOTA LLMs. All models improve with feedback.